# A Novel Hybrid Deep Learning Technique for Speech Emotion Detection using Feature Engineering

### v1

@@ 4.1 Performance Matrices and Evaluation
We also performed 5-fold cross-validation with PCA-reduced features (principle component analysis) to evaluate model consistency [38, 39]. As seen in Table 4, the results remained highly stable across datasets, confirming the robustness of DeepCRF. The close agreement between 80-20 split and cross-validation accuracies highlights the model’s generalization ability, while PCA successfully reduced feature size without loss of accuracy.
@@ 4.2 Comparison with the State of the Art
Although [23] achieved slightly higher performance on the Savee dataset (97.11%) and achieved the same performance on the combined dataset (99.18%), our proposed model remains highly competitive while demonstrating robustness across all datasets. Furthermore, there is no other model found in literature that worked with the combined five datasets, where our model achieved an accuracy of 93.76%.
Moreover, in comparison to ensemble-based architectures such as the 1D CNNs-LSTM-GRU model by [3], and ConvLSTM models [23], the proposed DCRF-BiLSTM architecture—combining deep contextualized BiLSTM layers with a CRF-based sequence labeling mechanism—demonstrates superior generalization for audio signals. This confirms the model’s ability to capture temporal dependencies and contextual emotional insights more effectively. The proposed model shows consistent performance for all datasets, ensuring the robustness and adaptability of different emotional speech corpora.
@@
S. Mekruksavanich[40] 75.83 55.71 65.83 - 65.77 - - DCNN
U. Asiya [41] 68.00 - - - - 89.00 - 1D CNN
D. Issa [24] 71.61 - - 86.1 - - - 1D CNN LSTM, 5-fold cross validation
L. T. C. Ottoni [16] 97.01 100.00 90.62 - 83.28 97.37 - 2 CNN+LSTM, Split (80-10-10)
S. Jothiman [18] 92.60 99.60 84.90 - 89.90 - - CNN + LSTM model using the 80%-20% train-test split
M. R. Ahmed [3] 95.62 99.46 93.22 95.42 74.09 - - 1D CNNs-LSTM-GRU-ENSEMBLE
R. M. Ben-Sauo [23] 96.83 99.99 97.11 98.65 - 99.18 - CONVLSTM model using the 5-fold-cross validation

### v2

@@ 4.1 Performance Matrices and Evaluation
We also performed 5-fold cross-validation with PCA-reduced features (principle component analysis) to evaluate model consistency [abdi2010principal, benba2017voice]. As seen in Table 4, the results remained highly stable across datasets, confirming the robustness of DeepCRF. The close agreement between 80-20 split and cross-validation accuracies highlights the model’s generalization ability, while PCA successfully reduced feature size without loss of accuracy.
@@ 4.2 Comparison with the State of the Art
Although [ben2024enhancing] achieved slightly higher performance on the Savee dataset (97.11%) and achieved the same performance on the combined dataset (99.18%), our proposed model remains highly competitive while demonstrating robustness across all datasets. Furthermore, there is no other model found in literature that worked with the combined five datasets, where our model achieved an accuracy of 93.76%.
Moreover, in comparison to ensemble-based architectures such as the 1D CNNs-LSTM-GRU model by [ahmed2023ensemble], and ConvLSTM models [ben2024enhancing], the proposed DCRF-BiLSTM architecture—combining deep contextualized BiLSTM layers with a CRF-based sequence labeling mechanism—demonstrates superior generalization for audio signals. This confirms the model’s ability to capture temporal dependencies and contextual emotional insights more effectively. The proposed model shows consistent performance for all datasets, ensuring the robustness and adaptability of different emotional speech corpora.
@@
S. Mekruksavanich[mekruksavanich2020negative] 75.83 55.71 65.83 - 65.77 - - DCNN
U. Asiya [asiya2021speech] 68.00 - - - - 89.00 - 1D CNN
D. Issa [issa2020speech] 71.61 - - 86.1 - - - 1D CNN LSTM, 5-fold cross validation
L. T. C. Ottoni [ottoni2023deep] 97.01 100.00 90.62 - 83.28 97.37 - 2 CNN+LSTM, Split (80-10-10)
S. Jothiman [jothimani2022mff] 92.60 99.60 84.90 - 89.90 - - CNN + LSTM model using the 80%-20% train-test split
M. R. Ahmed [ahmed2023ensemble] 95.62 99.46 93.22 95.42 74.09 - - 1D CNNs-LSTM-GRU-ENSEMBLE
R. M. Ben-Sauo [ben2024enhancing] 96.83 99.99 97.11 98.65 - 99.18 - CONVLSTM model using the 5-fold-cross validation
@@ 5 Conclusions
Declaration on Generative AI
The machine learning code in this research was implemented by the author(s), with limited assistance from ChatGPT for error checking and improvement. In addition, AI tools such as ChatGPT and Grammarly were used for grammar and spelling checks, paraphrasing, and rewording during the preparation of this manuscript. After using these tools, the author(s) reviewed and edited all content as needed and assume(s) full responsibility for the content of the publication.